Find the equation of the line passing through the point (7,2) and perpendicular to the line y=5x-6 Write your answer in simplified fraction form. Do not use mixed fractions.
Found 2 solutions by algebrahouse.com, stanbon:
Answer by algebrahouse.com(1659)

point (7,2) and perpendicular to y = 5x - 6

Perpendicular lines have slopes which are negative reciprocals

y = 5x - 6
slope = 5
perpendicular slope is -1/5

(7,2), m = -1/5
y - y1 = m(x - x1) {point-slope form}
y - 2 = (-1/5)(x - 7) {substituted point and slope into point-slope form}
y - 2 = (-1/5)x + 7/5 {used distributive property}
y = (-1/5)x + 17/5 {added 2, which is 10/5, to each side}

For more help from me, visit: www.algebrahouse.com


Answer by stanbon(75887)

Find the equation of the line passing through the point (7,2) and perpendicular to the line y=5x-6 Write your answer in simplified fraction form. Do not use mixed fractions.
----------
The given equation has slope = 5
---
All perpendicular lines has slope = -1/5
---
Form: y = mx + b
Solve for "b":
2 = (-1/5)7 + b
---------
10/5 = -7/5 + b
----
b = 17/5
-----------------
Equation:
y = (-1/5)x + (17/5)
========================
Cheers,
Stan H.
========================